You currently have 325 lb of a cement mixture containing 18.0% sand by mass.
How much sand must you add to create a mixture that is 25.0% sand by mass?

Answer by mananth(16946)

You currently have 325 lb of a cement mixture containing 18.0% sand by mass.
How much sand must you add to create a mixture that is 25.0% sand by mass?

You currently have 325 lb of a cement mixture containing 18.0% sand =58.5 lbs sand
Assume we add x lbs of sand
58.5+x = (325+x)*0.25
58.5+x = 81.25 +0,25x
x-0.25x = 81.25-58.5=22.75
0.75x=22.75
x= 22.75/0.75= 30.33 approx
Check
Total Sand / Total mix
(58.5+30.33)/(325+30.33) =0.2499 = 24.99 %